### v3

@@ 6.5.2 Reliability Analysis
Additionally, we calculated Cronbach’s Alpha if an item is deleted for each criterion. This metric helps assess the impact of removing an individual item on the scale’s overall reliability. As summarized in Table 5, the "Alpha if Deleted" values are consistently close to the overall Cronbach’s Alpha. Specifically, these values range from 0.9699 (Usefulness) to 0.9798 (Informativeness), with all values slightly lower than the overall result. This suggests that removing any single item would only have a minimal impact on the reliability of the scale, indicating that each criterion contributes positively to the measurement of overall quality.
Further, we examined the item-total correlations, which quantify the strength of the relationship between each criterion and the total score. These correlations range from 0.9286 for the Informativeness criteria to 0.9660 for Usefulness. The high item-total correlations support the relevance of each item in assessing the key dimensions of the robot’s explanations.
Table 5: Cronbach’s Alpha and item-wise "Alpha if Deleted" in human evaluation.
@@
Finally, we assessed the inter-item correlations, which provide insight into how strongly the criteria correlate with one another. The inter-item correlation matrix, presented in Table 6, shows values ranging from 0.878 (between Informativeness and Satisfaction) to 0.981 (between Satisfaction and Usefulness). These high correlations indicate that the criteria are closely related but not redundant, as each captures a unique and complementary aspect of the evaluation process.
Table 6: Inter-Item correlation matrix in human evaluation.
@@
6.5.3 Correlation Analysis
Building on the findings from the reliability analysis, we further explored the relationships among the variables in our dataset through Pearson correlation matrices. While Cronbach’s Alpha results demonstrated the internal consistency of the evaluation criteria, the correlation analysis extends this by examining how these criteria relate to each other across different scenarios and explanation categories.
Figure 14 presents the Pearson correlation matrix for participants’ responses across the three scenarios. It reveals strong associations, indicating a high level of consistency in participants’ assessments. This consistency suggests that the explanations provided were clear and easily comparable, allowing participants to reliably evaluate them regardless of the specific context of the robot’s task.
Figures 15 and 16 depict the relationships among evaluation criteria and question categories. Regarding the criteria, Completeness and Informativeness exhibit a strong positive correlation, as do Understandability and Usefulness. This indicates that participants are more likely to perceive explanations as useful when they are also clear. In contrast, the correlation between Informativeness and Usefulness is weaker (0.55), highlighting a divergence in participants’ evaluations. This divergence may stem from biases previously discussed, where participants find it challenging to judge the usefulness of an explanation that is not fully accurate.
Regarding the question categories, the ‘Navigation Process Overview’ and ‘Specifics About Goals and Locations’ categories exhibit a strong correlation (0.99), indicating that participants tended to provide similar ratings for explanations in these areas. This suggests that these types of explanations are perceived similarly. In contrast, the ‘Trajectory Planning and Re-planning’ category shows weak or negative correlations with other categories. This discrepancy arises from differences in evaluating broad, open-ended explanations versus more specific ones.
@@ 6.5.3 Correlation Analysis
Figure 16: Correlation matrix of responses by category.
6.5.4 Inferential Statistics
@@
Statistical analysis often requires understanding the underlying data distribution to ensure the appropriate methods are applied. Since parametric tests assume normality, violations of this assumption can compromise the validity of the results. To assess normality, we applied the Shapiro-Wilk test, a robust method for small to moderate-sized samples, across various groupings, including scenarios, categories, questions, and criteria.
@@
The test results consistently revealed significant deviations from normality, with $p$-values well below 0.05, leading to the rejection of the null hypothesis of normality for all groups. These findings confirm that the dataset does not satisfy the assumptions required for parametric tests, such as ANOVA or t-tests. As a result, non-parametric methods were adopted in subsequent analyses to ensure the robustness and validity of statistical inferences.
@@
The Friedman test was used due to its ability to assess differences in related groups when the data does not meet the assumptions of normality required for parametric tests. The results of the Friedman test across the three scenarios reveal significant differences in participant ratings for several explanations, indicating varying levels of agreement across conditions. In Scenario 1, significant differences were observed for explanations to Questions 1 and 2, suggesting diverse responses. In contrast, most other questions in this scenario showed no significant differences ($p\geq 0.05$).
@@
Explanations 1 and 2 in all three scenarios addressed broad questions, whose answers diverged from the ground truth, as discussed in the GPT-4 evaluation in 6.3. Broad questions often result in responses that prioritize varied information, such as BTs, locations, or timestamps, making it difficult to align the answers with the ground truth. This fact justifies the significant variability in participant ratings for these explanations.
@@
In Scenario 2, significant differences were observed for Explanations 1, 7, 8, and 15, with $p$-values below the 0.05 threshold, while other questions did not show significant variations. Specifically, the explanations for Questions 7 and 8 were misaligned with the ground truth, leading to significant variability in participant assessments. In these cases, participants’ ratings lacked a clear trend across the evaluated criteria. A similar pattern was observed in Explanation 15 of Scenario 2, where the model provided irrelevant information that did not directly address the user’s inquiry. This lack of relevance and alignment likely contributed to the observed variability in participant ratings.
@@
Finally, Scenario 3 showed significant differences for Explanations 1, 2, 7, and 16, indicating that the conditions for these explanations led to notable variations in participant ratings. Explanation 7, does not address the answer to the corresponding question, including no relevant information. Explanation 16 in Scenario 3 caused variability in participants’ answers due to its incomplete nature, highlighting the importance of content clarity in generating consistent and meaningful explanations. These findings suggest that while most explanations remain consistent, certain ones resonate differently depending on the scenario, nature of the question, and LLMs indeterminism.
@@
The post-hoc Nemenyi test was conducted for explanations that showed significant differences in participant ratings based on the Friedman test results. This allowed us to identify specific participant pairs with notable divergences in their responses across the three scenarios. The Nemenyi test revealed that most participants were consistent in their evaluations, but certain pairs exhibited significant variability. For example, in Scenario 1, the evaluation for Explanations 1 and 2 highlighted differences involving Participant 8 and several others. In Scenario 2, significant differences emerged for Explanation 1 among Participants 3, 6, 7, and 13. Similarly, in Scenario 3, variability was observed for Explanations 1 and 16, where selected participants differed significantly from others. These discrepancies are likely due to natural human biases, differences in interpretation, and subjective judgments that are inherent in human responses. Overall, while the explanations were generally clear and consistent for most participants, the observed variability underscores the role of individual perspectives and biases in shaping evaluations, reflecting the inherent complexity of human judgment.
6.6 Similarity Between Human and LLM Evaluations
@@
As described in 5.2, previous studies have demonstrated that LLMs can effectively approximate human judgment in various evaluation tasks, including those involving subjective interpretation. In this work, we conducted a focused validation of LLM-generated evaluations by comparing a subset of 48 explanations from one of six evaluation runs performed by the LLM against human assessments. This subset included 16 explanations per scenario, allowing us to validate the LLM’s judgment within a practical scope.
@@
Figures 18 and 19 depict the mean responses from 17 human participants and GPT-4, respectively, acting as a judge in this subset. The comparison highlights notable similarities in scoring patterns across scenarios and explanations, suggesting a high degree of alignment between human and LLM evaluations. Both evaluations consistently identify high-quality explanations with scores clustered around 4.5–5, indicating shared agreement on the quality of certain explanations. Similarly, lower-rated explanations identified by humans are also flagged by the LLM, revealing a mutual recognition of weaker content.
@@
However, while the overall trends are similar, there are some differences in the exact ratings and outlier values between the two evaluations. These differences could be attributed to nuances in how the LLM interprets the explanations compared to human evaluators, who might take into account subtle contextual or subjective factors that an LLM might miss. Despite these slight variations, the visual correspondence between the human and LLM heatmaps indicates that the LLM provides a reliable approximation of human judgment, making it a promising tool for scaling evaluation processes without sacrificing much in terms of accuracy. This similarity could be valuable in demonstrating that LLMs are suitable alternatives for tasks like explanation evaluation, especially in large-scale studies where human evaluation may be impractical.
@@
Figure 18: Heatmap of mean responses from human evaluators Across Scenarios and Explanations.
Figure 19: Heatmap of mean responses from the LLM-as-Judge evaluation across scenarios and explanations.
7 Conclusions
This work presents an accountability and explicability architecture, implemented for Robot Operating System (ROS)-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, providing satisfactory results for users and thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging Application Programming Interface (API), easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
The results achieved highlight the capability of our solution to guarantee the immutability of information generated by an agent without compromising its performance, a feature particularly important when trying to identify the causes behind an event. Moreover, our system improves efficiency and effectiveness in the deployment of autonomous agents, providing natural language explanations across key navigation aspects including goal achievement, obstacle appearance, trajectory re-planning, and objective cancellation or abortion. This fact strengthens trust and confidence in the use of autonomous agents across diverse applications and contexts.
@@ 7 Conclusions
While our approach has shown promising results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use. Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects. Therefore, future work should explore the impact of In-Context Learning techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context. This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating Knowledge Graphs with vector indexes. Furthermore, by incorporating real-time explanation generation mechanisms into our solution, the responsiveness and adaptability of the system could be improved. These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
@@ A.2 Explanations Obtained
. Additionally, a Python notebook has been developed to interact with the content of the previous sheet. This notebook at the following link  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@
[…]
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
Data and tools are available in a public repository  [^6] https://github.com/laurafbec/immutable_explainable_BBR.
Acknowledgments

### v4

@@ VI-E2 Reliability Analysis
Additionally, we calculated Cronbach’s Alpha if an item is deleted for each criterion. This metric helps to assess the impact of removing an individual item on the scale’s overall reliability. As summarized in Table V, the ”Alpha if Deleted” values are consistently close to the overall Cronbach’s Alpha. Specifically, these values range from 0.9699 (Usefulness) to 0.9798 (Informativeness), with all values being slightly lower than the overall result. This suggests that removing any single item would have only a minimal impact on the reliability of the scale, indicating that each criterion contributes positively to the measurement of overall quality.
Further, we examined item-total correlations, which quantify the strength of the relationship between each criterion and the total score. These correlations range from 0.9286 for the Informativeness criteria to 0.9660 for Usefulness. The high item-total correlations support the relevance of each item in assessing the key dimensions of the robot’s explanations.
TABLE V: Cronbach’s Alpha and item-wise ”Alpha if Deleted” in human evaluation.
@@
Finally, we assessed inter-item correlations, which provide insight into how strongly the criteria correlate with one another. The inter-item correlation matrix, presented in Table VI, shows values ranging from 0.878 (between Informativeness and Satisfaction) to 0.981 (between Satisfaction and Usefulness). These high correlations indicate that the criteria are closely related but not redundant, as each captures a unique and complementary aspect of the evaluation process.
TABLE VI: Inter-Item correlation matrix in human evaluation.
@@
VI-E3 Correlation Analysis
Building on the findings of the reliability analysis, we further explored the relationships among the variables in our dataset using Pearson correlation matrices. While Cronbach’s Alpha results demonstrated the internal consistency of the evaluation criteria, the correlation analysis extends it by examining how these criteria relate to each other across different scenarios and explanation categories.
Figure 14 presents the Pearson correlation matrix for participants’ responses across the three scenarios. This reveals strong associations, indicating a high level of consistency in the participants’ assessments. This consistency suggests that the explanations provided were clear and easily comparable, allowing participants to reliably evaluate them regardless of the specific context of the robot’s task.
Figure 15 depicts the relationships among the evaluation criteria. Completeness and Informativeness exhibit a strong positive correlation, as do Understandability and Usefulness. This indicates that participants are more likely to perceive explanations as useful when they are clear. In contrast, the correlation between Informativeness and Usefulness is weaker (0.55), highlighting a divergence in participants’ evaluations. This divergence may stem from the biases previously discussed, where participants find it challenging to judge the usefulness of an explanation that is not fully accurate.
@@ VI-E3 Correlation Analysis
VI-E4 Inferential Statistics
@@
Statistical analysis often requires an understanding of the underlying data distribution to ensure that appropriate methods are applied. As parametric tests assume normality, violations of this assumption can compromise the validity of the results. To assess normality, we applied the Shapiro-Wilk test, a robust method for small to moderate-sized samples, across various groupings, including scenarios, categories, questions, and criteria.
@@
The test results consistently revealed significant deviations from normality, with $p$-values well below 0.05, leading to rejection of the null hypothesis of normality for all groups. These findings confirm that the dataset does not satisfy the assumptions required for parametric tests, such as ANOVA or t-tests. As a result, non-parametric methods were adopted in subsequent analyses to ensure the robustness and validity of statistical inferences.
@@
The Friedman test was used because of its ability to assess differences in related groups when the data does not meet the assumptions of normality required for the parametric tests. The results of the Friedman test across the three scenarios reveal significant differences in participant ratings for several explanations, indicating varying levels of agreement across conditions. In Scenario 1, significant differences were observed for the explanations of Questions 1 and 2, suggesting diverse responses. In contrast, most of the other questions in this scenario showed no significant differences ($p\geq 0.05$).
@@
Explanations 1 and 2 in all three scenarios addressed broad questions, whose answers diverged from the ground truth, as discussed in the GPT-4 evaluation in VI-C. Broad questions often result in responses that prioritize varied information, such as BTs, locations, or timestamps, making it difficult to align the answers with the ground truth. This fact justifies the significant variability in the participant ratings for these explanations.
@@
In Scenario 2, significant differences were observed for Explanations 1, 7, 8, and 15, with $p$-values below the threshold of 0.05, while the other questions did not show significant variations. Specifically, the explanations for Questions 7 and 8 were misaligned with the ground truth, leading to significant variability in the participant assessments. In these cases, the participants’ ratings lacked a clear trend across the evaluated criteria. A similar pattern was observed in Explanation 15 of Scenario 2, where the model provided irrelevant information that did not directly address the user’s inquiry. This lack of relevance and alignment likely contributed to the observed variability in the participant ratings.
@@
Finally, Scenario 3 showed significant differences for Explanations 1, 2, 7, and 16, indicating that the conditions for these explanations led to notable variations in the participant ratings. Explanation 7, does not address the answer to the corresponding question, including no relevant information. Explanation 16 in Scenario 3 caused variability in participants’ answers due to its incomplete nature, highlighting the importance of content clarity in generating consistent and meaningful explanations. These findings suggest that while most explanations remain consistent, certain ones resonate differently depending on the scenario, nature of the question, and LLMs indeterminism.
@@
A post-hoc Nemenyi test was conducted for explanations that showed significant differences in participant ratings based on the Friedman test results. This allowed us to identify specific participant pairs with notable divergences in their responses across the three scenarios. The Nemenyi test revealed that most participants were consistent in their evaluations however, certain pairs exhibited significant variability. For example, in Scenario 1, the evaluation for Explanations 1 and 2 highlighted differences involving Participant 8 and several others. In Scenario 2, significant differences emerged for Explanation 1 between Participants 3, 6, 7, and 13. Similarly, in Scenario 3, variability was observed for Explanations 1 and 16, where the selected participants differed significantly from the others. These discrepancies are likely due to natural human biases, differences in interpretation, and subjective judgments inherent in human responses. Overall, while the explanations were generally clear and consistent for most participants, the observed variability underscores the role of individual perspectives and biases in shaping evaluations, reflecting the inherent complexity of human judgment.
VI-F Similarity Between Human and LLM Evaluations
@@
As described in V-B, previous studies have demonstrated that LLMs can effectively approximate human judgment in various evaluation tasks, including those involving subjective interpretation. In this study, we conducted a focused validation of LLM-generated evaluations by comparing a subset of 48 explanations from one of the six evaluation runs performed by the LLM against human assessments. This subset included 16 explanations per scenario, allowing us to validate the LLM judgment within a practical scope.
@@
Figures 17(a) and 17(b) depict the mean responses from the 17 human participants and GPT-4, respectively, acting as a judge in this subset. The comparison highlights notable similarities in scoring patterns across scenarios and explanations, suggesting a high degree of alignment between human and LLM evaluations. Both evaluations consistently identify high-quality explanations with scores clustered around 4.5–5, indicating a shared agreement on the quality of certain explanations. Similarly, lower-rated explanations identified by humans are also flagged by the LLM, revealing mutual recognition of weaker content.
@@
However, although the overall trends are similar, there are some differences in the exact ratings and outlier values between the two evaluations. These differences could be attributed to nuances in how the LLM interprets the explanations compared to human evaluators, who might consider subtle contextual or subjective factors that an LLM might miss. Despite these slight variations, the visual correspondence between the human and LLM heatmaps indicates that the LLM provides a reliable approximation of human judgment, making it a promising tool for scaling evaluation processes without sacrificing much in terms of accuracy. This similarity could be valuable in demonstrating that LLMs are suitable alternatives for tasks like explanation evaluation, especially in large-scale studies where human evaluation may be impractical.
(a) Human evaluation.
(b) LLM-as-Judge evaluation.
@@
Figure 17: Heatmaps of mean responses across scenarios and explanations. (a) Human evaluators. (b) LLM-as-Judge evaluation.
VII Conclusions
This work presents an accountability and explainability architecture for Robot Operating System (ROS)-based mobile robots. The proposed system comprises a black box component that ensures the verifiability and integrity of logged messages using blockchain technology, and an explainability component based on Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs). This private question-answering system leverages the data recorded by the accountability component to provide context for generating natural language explanations. The results demonstrate the feasibility of producing accurate, coherent, concise, and helpful explanations from raw Rosbag data, thereby supporting user understanding and enhancing human-robot interactions.
Both components are integrated through a third main module in charge of filtering non-relevant information and transforming technical content into more accessible messages. This functionality reduces the complexity for developers by eliminating the need for direct interaction with the logging Application Programming Interface, thereby easing the development process and improving the adaptability and scalability of the system across diverse navigation tasks.
Our evaluation confirms that the system guarantees data immutability without compromising performance, supporting post-hoc analysis and failure investigation. Specifically, hashing one of every 100 messages resulted in less than 5% message loss, even for high-frequency topics exceeding 300 Hz, with minimal resource consumption. These results validate the efficiency and practicality of accountability mechanisms.
The system also provides effective explanations across key aspects of autonomous navigation, such as goal completion, obstacle detection, trajectory re-planning, and task cancellation or abortion. The generated explanations achieved high scores for accuracy, coherence, and user satisfaction. LLM-as-a-judge evaluations reported correctness rates at or above 75% across all scenarios, while human evaluations showed strong agreement, with a Cronbach’s Alpha of 0.9792, indicating high internal consistency. The alignment between human and LLM judgments suggests that LLMs can serve as reliable and scalable alternatives to human evaluation for assessing explanation quality.
These findings highlight the potential of the proposed architecture to support the broader deployment of autonomous agents in safety-critical scenarios. By enabling reliable post-hoc inspection and producing accessible, high-quality explanations, our system bridges the gap between complex robotic behavior and human interpretation. This capability is particularly relevant in contexts that require transparency, positioning our approach as a practical foundation for responsible and explainable robots.
@@ VII Conclusions
While our solution has shown promising results in navigation functionalities across various scenarios, future work should extend its application to other domains by tailoring the recording and processing components to specific topics relevant to each context. Furthermore, the consistency and robustness of the responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects. To this end, future efforts should explore the impact of In-Context Learning techniques on prompt formulation, along with fine-tuning the model to better align with the unique demands of autonomous agent tasks.
Additionally, our RAG approach could be further optimized to retrieve more relevant information from the available context. This improvement may involve structuring the vector database into separate collections tailored to specific information types or adopting a hybrid retrieval strategy that combines Knowledge Graphs with vector-based indexing. Incorporating real-time explanation generation into the architecture could also enhance the responsiveness and adaptability of the system. Together, these enhancements have the potential to significantly improve the accuracy, relevance, and consistency of the generated explanations across diverse scenarios, thereby strengthening the overall effectiveness of the proposed architecture.
@@ A-B Explanations Obtained
. A Python notebook has been developed to interact with the content of the previous sheet. This notebook is available at the following link  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@
Acknowledgment